OpenStudy (anonymous):

Suppose that both 75% brass and 90% brass are available. How many pounds of each must be mixed to get 150 pounds of 87% brass? My book answers were x=30 pounds of 75% brass and y= 120 pounds of 90% brass. I wanted to know how to get that answer-- what is the equation is it x + y= 150 and ???????

OpenStudy (anonymous):

30*0.75+120*0.9=150*0.87

OpenStudy (anonymous):

so it is x*0.75+y*0.9=150*0.87

OpenStudy (anonymous):

plus x+y=150

OpenStudy (anonymous):

and if you solve these two equations than you get the result

OpenStudy (anonymous):

I will key in with my scientific calculator. sounds right wait.... thank you

OpenStudy (anonymous):

but the 30 and 120 is not given..... that is the answer at the back of my book

OpenStudy (anonymous):

yes that is what confused me :D

OpenStudy (anonymous):

hey he gave us a clue that brass is made of copper and zinc--- it means the other percentage to make a 100% is a zinc.

OpenStudy (anonymous):

but now it is clear. let x be the 75% one and y the 90% than we want to make 150 pounds so x+y=150 also we want to get 87% brass so x*0.75+y*0.9=150*0.87

OpenStudy (anonymous):

y=150-x from first equation you plug this into the second 0.75x +(150-x)*0.9=150*0.87

OpenStudy (anonymous):

this gives 4.5=0.15x so x=30 y=(150-x)=120

OpenStudy (anonymous):

I am using scientific calculator using the command matrix.... ok let me do that.

OpenStudy (anonymous):

calculators are bad... they do the work for you but you dont think

OpenStudy (anonymous):

hehehe! yah.... but we are allowed to use it though in math modeling... i hate math ill be glad to get out in this class... its the only math i will be taking in nursing. thank goodness

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

did you get the result by the way?

OpenStudy (anonymous):

not yet im still trying to figure out how to key in figures using matrix command.

OpenStudy (anonymous):

I did not use a calculator and had the answer 10 minutes ago :-)

OpenStudy (anonymous):

hahahaha!

OpenStudy (anonymous):

thanks for your time andras--- I got the answer solving manually using your equation. But i donno how to key in the scientific calculator though--thanks for your effort. love this site.